# MythTriage: Scalable Detection of Opioid Use Disorder Myths on a Video-Sharing Platform

### v1

@@ 8 Conclusion & Future Work
Applications for Public Health.MythTriage can empower public health practitioners monitor real-time misinformation trends, identify emerging myths, and launch targeted interventions promptly. Such insights can help clinicians understand common myths patients may encounter online, enabling better-informed, trust-building interactions.
Platform Moderation and Auditing.MythTriage can be integrated into platform moderation workflows to flag high-risk content and support scalable content auditing. Third-party researchers can also use the method to evaluate how various algorithms and filter settings shape user exposure to misinformation at scale, informing platform transparency and algorithmic accountability.
@@ 9 Limitations
Expanding Myths and Topics. In this work, we examined 8 expert-validated OUD-related myths and 8 topics, consisting of 4 treatment and 4 opioid topics (§3.1). While this covers more ground than prior works Mittal et al. (2024); ElSherief et al. (2021b), many other myths, such as “It Is Expensive To Treat Patients With OUD” AMCP (2023), and topics (e.g., Oxycontin) remains underexplored. Future works can extend MythTriage and our annotation guidelines to cover a broader range of opioid-related myths and topics.
Exploring Deeper and Beyond YouTube. This work presents the first large-scale analysis of OUD-related myths on YouTube. While prior research highlights the role of personalization in amplifying problematic content Hussein et al. (2020), future work can examine how personalization—such as user demographics, search history, and prior engagement— shapes exposure to OUD-related myth. Other platforms like BitChute Nguyen et al. (2024) and TikTok Boeker and Urman (2022) also merit investigation. Extending this analysis across platforms would enable cross-platform comparisons of myth prevalence. Additionally, while we can analyze what happens when a hypothetical user interacts on the platform, future work should incorporate crowdsourced data collection Juneja et al. (2023) to understand how real users encounter and engage with OUD-related misinformation.
@@
More Extensive Data Collection. In our work, we collected 2.9K search results and 343K recommendations, providing the first large-scale analyses on the OUD myth prevalence on YouTube. Future works can expand this by collecting data over longer periods to capture longitudinal trends and better understand how myths evolve over time.
Model Misclassification. We employ LLMs and distill lightweight classifiers validated on clinical expert-labeled datasets. Despite extensive model training and experimentation, the error rates in our MythTriage pipeline may potentially influence our downstream analysis. While our text-based approach using models like DeBERTA-v3 balances performance and efficiency (§5.3), future work can improve performance by incorporating multimodal features (e.g., thumbnails, video frames) Abdali et al. (2024), though this may increase cost, and integrating additional deferral mechanisms between experts and LLMs using uncertainty metrics from Farr et al. (2024a). MythTriage can be adapted to use small LLMs such as Llama-3-8B Zhan et al. (2025) for potentially better performance.
@@ 10 Ethical Considerations
To minimize exposure to harmful content, we did not involve crowd workers and instead worked with clinical researchers in substance use. Following Kirk et al. (2022), we conducted regular check-ins and debriefs to safeguard all data handlers. All analyzed videos were publicly available at the time of collection, and we only used video metadata (titles, descriptions, transcripts), without accessing user-specific data. In line with best practices Proferes et al. (2021), we will release only video IDs and labels to ensure reproducibility while protecting user privacy. Practitioners can use these IDs to retrieve their metadata via YouTube Data API.
@@ Appendix A Details on Curating Topics and Queries
Given our focus on detecting OUD-related misinformation, we selected 8 keywords as topics—the top four opioid and top four treatment keywords—capturing the most popular topics across both opioids and medication-assisted treatment. These opioid and MAT-related keywords are both widely-searched in the context of opioid use and treatment, which are often associated with OUD-related myths on social media platformsElSherief et al. (2024a).
@@ B.3 Annotation Process
To construct the gold-standard dataset, six clinical researchers, as experts, annotated 310 videos over three rounds of annotations. The clinical researchers included undergraduate and graduate research assistants as well as full-time professionals, all based at a U.S. university medical school. We did not provide payment, but we obtained their consent to use their annotations for our study.
In the first round, all six experts independently annotated 20 videos, familiarizing themselves with the annotation guidelines (Appendix B.2), providing feedback, and discussing disagreements. In the second round, they split into two groups of three annotators, annotating 40 videos per group. In the third round, they formed three pairs (e.g., groups of two), annotating 70 videos per group. Each round concluded with a debrief with the experts. Thus, in total, 210 videos had 2 annotators, 80 videos had 3 expert annotators, and 20 videos had 6 expert annotators.
@@ C.3 LLM Evaluation Results
Using both zero-shot and few-shot prompts, we evaluated 10 LLMs on the gold-standard dataset across 8 OUD-related myths. Detailed performance results of all 10 LLMs are shown in Tables 19, 20, and 21, with a summary of the best macro F1-scores by model in Table 12.
@@
Table 12: The best model performances (Macro F1-Score) across 8 OUD-related myths. For each myth, we bolded the best model performance. For each myth, we evaluate the performance on 305 videos from the expert-annotated gold standard dataset, excluding the five few-shot examples used in the prompt. Note that GPT-4o: GPT-4o-2024-08-06, GPT-4o-mini: GPT-4o-mini-2024-07-18, Claude-Sonnet: Claude-3.5-Sonnet-20241022, Claude-Haiku: Claude-3.5-Haiku-20241022, Llama-3-8B: Meta-Llama-3-8B-Instruct, Llama-3.3-70B: Meta-Llama-3.3-70B-Instruct, DeepSeek: DeepSeek-v3, Gemini-Pro: Gemini-1.5-Pro, Gemini-Flash: Gemini-1.5-Flash, Qwen-72B: Qwen-2.5-72b-instruct. Please refer to Tables 19-21 for the full performance metrics using both zero-shot and few-shot prompts.
@@ E.3 Results
We apply each deferral method to the classification of Myth 1 on the gold-standard dataset to evaluate their performance. In Table 15, MC-Dropout and Softmax Entropy achieve the highest macro F1-score (0.87), but they do so by deferring 90% of predictions to GPT-4o. While this boosts performance, such high deferral rates severely undermine the purpose of MythTriage—effectively shifting most of the world to the expensive LLM and compromising scalability and cost-efficiency. Additionally, MC-Dropout is computationally intensive, requiring multiple forward passes per input. This further increases computational cost and latency, making it an unattractive option despite its predictive performance.
@@ F.2 Approach 2: GPT-4o
  • Financial Cost: Each prompt used on average 6,066.92 input tokens and 144.01 output tokens. Using OpenAI API pricing OpenAI (2025), each myth-level prompt would roughly costs (6,066.92 input tokens ×\$$\times\frac{\textrm{\$2.50}}{\textrm{1M input tokens}}$) + (144.01 output tokens ×\$$\times\frac{\textrm{\$10.00}}{\textrm{1M output tokens}})$ = \$0.0166. For labeling 8 myths across all 164K videos, it is estimated to cost: $164{,}085\times 8\times\$0.0166$ = \$21,790.49.
@@ F.3 Approach 3: MythTriage
[…]
@@ I.1 Myth Bias Scores in Search Queries
Figure 7 displays the top 15 search queries with the highest myth bias scores, indicating higher prevalence of myths. The query "methadone detox" has the highest bias score, implicitly reinforcing three myths: that methadone is as dangerous or addictive as opioids (Myth 1), that the ultimate goal of MAT is abstinence from any opioid use (Myth 3), and that detoxing is a safe and valid treatment approach (Myth 6). Notably, even seemingly neutral queries such as "methadone" and "kratom" yield search results biased toward misinformation, highlighting the pervasive influence of myth-supporting content in OUD-related search results.
@@ I.3 Distribution of Labels Across Myths and Filters
Figure 9 shows the frequency of labels across the eight myths and overall stance, grouped by filters. Sorting by relevance shows high counts of opposing labels across all myths, as discussed in §7.1. However, employing a different search filter returns relatively less opposing labels and more supporting labels.
@@ I.4 Engagement Analysis
A Shapiro-Wilk Test revealed that our data is not normal, so we used the non-parametric Kruskal-Wallis test to compare video engagement across stance labels (e.g., supporting, neither, and opposing). We found significant differences in views (KW H(2, N$=$1776)=10.93, p$<$0.01, $\eta^{2}$=0.005) and number of comments (KW H(2, N$=$1776)=12.38, p$<$0.01, $\eta^{2}$=0.0.006), but not in likes and like-to-view ratio. We conducted post-hoc Conover-Iman Test with Bonferroni adjustment, revealing that opposing videos tend to get more views than supporting and neither videos. Meanwhile, we found that supporting videos tend to get more comments than neither videos, indicating higher engagement and discussion around myth-supporting content.
@@
VIDEO__\__TAGS: [EXAMPLE1_VIDEO_TAGS]

### v2

@@ 8 Conclusion & Future Work
Applications for Public Health.MythTriage can empower public health practitioners to monitor real-time misinformation trends, identify emerging myths, and launch targeted interventions promptly. Such insights can help clinicians understand common myths patients may encounter online, enabling better-informed, trust-building interactions.
Platform Moderation and Auditing. Video-sharing platforms like YouTube have introduced policies to address harmful and unsubstantiated content Halprin and O’Conner (2022), including medical misinformation policies targeting treatment-related claims Google (2025), indicating that misinformation mitigation is a platform priority. MythTriage can be integrated into platform moderation workflows to flag high-risk content, support scalable content auditing, and provide actionable insights that can inform at-scale moderation improvements. Third-party researchers can use MythTriage to evaluate how various algorithms and filter settings shape user exposure to misinformation at scale, informing platform transparency and algorithmic accountability.
@@ 9 Limitations
Expanding Myths and Topics. In this work, we examined 8 expert-validated OUD-related myths and 8 topics, consisting of 4 treatment and 4 opioid topics (§3.1). While this covers more ground than prior works Mittal et al. (2024); ElSherief et al. (2021), many other myths, such as “It Is Expensive To Treat Patients With OUD” AMCP (2023), and topics (e.g., OxyContin) remain underexplored. Future works can extend MythTriage and our annotation guidelines to cover a broader range of opioid-related myths and topics.
Exploring Deeper and Beyond YouTube. This work presents the first large-scale analysis of OUD-related myths on YouTube. While prior research highlights the role of personalization in amplifying problematic content Hussein et al. (2020), future work can examine how personalization—such as user demographics, search history, and prior engagement— shapes exposure to OUD-related myth. Other video-sharing platforms like BitChute Nguyen et al. (2024) and TikTok Boeker and Urman (2022) also merit investigation. Extending this analysis across platforms would enable cross-platform comparisons of myth prevalence.
Google Trends Alignment. To curate search topics and queries, we use Google Trends, configured to reflect YouTube Search, to capture highly popular opioid-related queries based on real-world search interests. This ensures our analysis reflects the amount of misinformation many YouTube users are likely to encounter when seeking OUD information, but may bias the dataset toward mainstream searches and underrepresent niche queries, including those used by stigmatized or vulnerable populations. Future work could address this gap through crowdsourced audit Juneja et al. (2023) or privacy-preserving audits to capture how real users, including those from vulnerable populations, encounter and engage with OUD-related misinformation.
@@
More Extensive Data Collection. In our work, we collected 2.9K search results and 343K recommendations, providing the first large-scale analyses on the OUD myth prevalence on YouTube. Future work can expand this by collecting data over longer periods to capture longitudinal trends and better understand how myths evolve over time.
Model Misclassification. We employ LLMs and distill lightweight classifiers validated on clinical expert-labeled datasets. Despite extensive model training and experimentation, the error rates in our MythTriage pipeline may potentially influence the distillation of the lightweight classifiers and our downstream analysis. While our text-based approach using models like DeBERTA-v3 balances performance and efficiency (§5.3), future work can improve performance by incorporating multimodal features (e.g., thumbnails, video frames) Abdali et al. (2024), though this would increase computational and financial costs of the pipeline by requiring multimodal LLMs, and integrating additional deferral mechanisms between experts and LLMs using uncertainty metrics from Farr et al. (2024a). MythTriage can be adapted to use small, open-weight LLMs such as Llama-3-8B Zhan et al. (2025) for potentially better performance, though even models with $\geq$ 1B parameters would substantially increase training and inference costs.
@@ 10 Ethical Considerations
To minimize exposure to harmful content, we did not involve crowd workers and instead worked with clinical researchers in substance use. Following Kirk et al. (2022), we conducted regular check-ins and debriefs to safeguard all data handlers. All analyzed videos were publicly available at the time of collection, and we only used video metadata (titles, descriptions, transcripts), without accessing user-specific data. In line with best practices Proferes et al. (2021), we will release only video IDs and labels to ensure reproducibility while protecting user privacy. Practitioners can use these IDs with the YouTube Data API to retrieve metadata and hydrate the released dataset.
@@
Acknowledgements
This paper was supported by the National Institute of Health’s NIDA grant DA056725-01A1 and the National Science Foundation grant CNS-2344925. We thank the Cavazos Lab at the Washington University School of Medicine for providing clinical research expertise and annotations in this study.
@@ Appendix A Details on Curating Topics and Queries
Given our focus on detecting OUD-related misinformation, we selected 8 keywords as topics—the top four opioid and top four treatment keywords—capturing the most popular topics across both opioids and medication-assisted treatment. These opioid and MAT-related keywords are both widely-searched in the context of opioid use and treatment, which are often associated with OUD-related myths on social media platformsElSherief et al. (2024).
@@ B.3 Annotation Process
To construct the gold-standard dataset, six clinical researchers, as experts, annotated 310 videos over three rounds of annotations. The clinical experts constituted a mixture of one clinical research scientist with a PhD, four graduate-level clinical researchers with Master’s degrees, and one undergraduate research assistant based at a flagship medical school in the U.S. These clinical researchers, including an undergraduate research assistant (with >1>1 year of research experience), regularly conduct clinical trials and perform research on substance use disorder and recovery, providing strong subject matter expertise. We did not provide payment, but we obtained their consent to use their annotations for our study.
In the first round, all six experts independently annotated 20 videos, familiarizing themselves with the annotation guidelines (Appendix B.2), providing feedback, and discussing disagreements. In the second round, they split into two groups of three annotators, annotating 40 videos per group. In the third round, they formed three pairs (e.g., groups of two), annotating 70 videos per group. Each round concluded with a debrief with the experts. Thus, in total, 210 videos had 2 annotators, 80 videos had 3 expert annotators, and 20 videos had 6 expert annotators. The annotation process spanned nearly three months and required close collaboration with clinical experts, who balanced these tasks alongside their own research.
@@ Appendix C Additional Details on LLM-Based Myth Detection
Prior work Jung et al. (2025); Juneja et al. (2023) has consistently shown that using all textual metadata–titles, descriptions, transcripts, and tags–yields the best performance for misinformation detection. For example, Jung et al. (2025) conducted ablation studies and found that removing any metadata component reduced performance in detecting COVID-19 misinformation in YouTube videos, indicating that each component contributes a useful signal. Based on these findings, we used the full set of text metadata components without re-running similar ablations, as our focus was on triage performance.
@@ C.3 LLM Evaluation Results
Using both zero-shot and few-shot prompts, we evaluated 10 LLMs on the gold-standard dataset across 8 OUD-related myths. Detailed performance results of all 10 LLMs are shown in Tables 20, 21, and 22, with a summary of the best macro F1-scores by model in Table 13.
@@
Table 13: The best model performances (Macro F1-Score) across 8 OUD-related myths. For each myth, we bolded the best model performance. For each myth, we evaluate the performance on 305 videos from the expert-annotated gold standard dataset, excluding the five few-shot examples used in the prompt. Note that GPT-4o: GPT-4o-2024-08-06, GPT-4o-mini: GPT-4o-mini-2024-07-18, Claude-Sonnet: Claude-3.5-Sonnet-20241022, Claude-Haiku: Claude-3.5-Haiku-20241022, Llama-3-8B: Meta-Llama-3-8B-Instruct, Llama-3.3-70B: Meta-Llama-3.3-70B-Instruct, DeepSeek: DeepSeek-v3, Gemini-Pro: Gemini-1.5-Pro, Gemini-Flash: Gemini-1.5-Flash, Qwen-72B: Qwen-2.5-72b-instruct. Please refer to Tables 20-22 for the full performance metrics using both zero-shot and few-shot prompts.
@@ E.3 Results
We apply each deferral method to the classification of Myth 1 on the gold-standard dataset to evaluate their performance. In Table 16, MC-Dropout and Softmax Entropy achieve the highest macro F1-score (0.87), but they do so by deferring 90% of predictions to GPT-4o. While this boosts performance, such high deferral rates severely undermine the purpose of MythTriage—effectively shifting most of the work to the expensive LLM and compromising scalability and cost-efficiency. Additionally, MC-Dropout is computationally intensive, requiring multiple forward passes per input. This further increases computational cost and latency, making it an unattractive option despite its predictive performance.
@@ F.2 Approach 2: GPT-4o
  • Financial Cost: Each prompt used on average 6,066.92 input tokens and 144.01 output tokens. Using OpenAI API pricing OpenAI (2025), each myth-level prompt would roughly costs (6,066.92 input tokens $\times\frac{\textrm{\$2.50}}{\textrm{1M input tokens}}$) + (144.01 output tokens $\times\frac{\textrm{\$10.00}}{\textrm{1M output tokens}})$ = \$0.0166. For labeling 8 myths across all 164K videos, it is estimated to cost: $164{,}085\times 8\times\$0.0166$ = \$21,790.49.
@@ F.3 Approach 3: MythTriage
    At 0.374 kg CO2/kWh U.S. Environmental Protection Agency (2025), the estimated emissions are = $282.15\times 0.374=\textbf{105.52 kg CO\textsubscript{2}}$.
@@ I.1 Myth Bias Scores in Search Queries
Figure 7 displays the top 15 search queries with the highest myth bias scores, indicating a higher prevalence of myths. The query "methadone detox" has the highest bias score, implicitly reinforcing three myths: that methadone is as dangerous or addictive as opioids (Myth 1), that the ultimate goal of MAT is abstinence from any opioid use (Myth 3), and that detoxing is a safe and valid treatment approach (Myth 6). Notably, even seemingly neutral queries such as "methadone" and "kratom" yield search results biased toward misinformation, highlighting the pervasive influence of myth-supporting content in OUD-related search results.
@@ I.3 Distribution of Labels Across Myths and Filters
Figure 9 shows the frequency of labels across the eight myths and overall stance, grouped by filters. Sorting by relevance shows high counts of opposing labels across all myths, as discussed in §7.1. However, employing a different search filter returns relatively fewer opposing labels and more supporting labels.
@@ I.4 Engagement Analysis
A Shapiro-Wilk Test revealed that our data is not normal, so we used the non-parametric Kruskal-Wallis test to compare video engagement across stance labels (e.g., supporting, neither, and opposing). We found significant differences in views (KW H(2, N==1776)=10.93, p<<0.01, $\eta^{2}$=0.005) and number of comments (KW H(2, N==1776)=12.38, p<<0.01, $\eta^{2}$=0.006), but not in likes and like-to-view ratio. We conducted post-hoc Conover-Iman Test with Bonferroni adjustment, revealing that opposing videos tend to get more views than supporting and neither videos. Meanwhile, we found that supporting videos tend to get more comments than neither videos, indicating higher engagement and discussion around myth-supporting content.
@@
VIDEO_\_TAGS: [EXAMPLE1_VIDEO_TAGS]